# Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective

### v1

@@ Abstract
Conclusion and Outlook.— In this Letter, we have investigated the distribution of Yang-Lee zeros in quantum phase transitions of the SSH model and the XXZ model and shown how Yang-Lee zeros enter the phase transition points. Especially, in the XXZ model, we have proposed the simplified Bethe ansatz equation (SBAE) to exactly solve the energy levels around the ground state and calculated the partition function in the zero-temperature limit. We have also explored the critical behavior at the edges of Yang-Lee zeros of both models. Furthermore, we find that the edges of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground state and discuss the mechanism behind the correspondence. We have also generalized the correspondence to an arbitrary PT-symmetric free-fermionic system.
@@
The Yang-Lee zeros in quantum phase transitions studied in this Letter are not only an interesting mathematical concept but also experimentally realizable. The non-Hermitian SSH model can be realized by dilating it into a larger Hermitian Hamiltonian due to the PT symmetry Zhang et al. (2021); Yu et al. (2022); Zeuner et al. (2015). We can introduce an ancillary qubit as an environment coupled to the targeted spin to realize the Hamiltonian (2). For the non-Hermitian XXZ model, we can realize it in a non-Hermitian Bose-Hubbard model in a strongly correlated regime Yamamoto et al. (2022); Xu and Chen (2020); Tomita et al. (2017), where the non-Hermiticity originates from the two-body loss in the open quantum systems. By coupling the target system to an ancilla, we expect the Yang-Lee zeros and its associated edge singularity to be measurable with the decoherence of the spin in the ancilla Peng et al. (2015); Wei (2017, 2018); Matsumoto et al. (2022). The Yang-Lee zeros in those systems can be also measured from the diagonal elements of the time-evolution operator $U=\exp(i\tilde{H}t)$, where $\tilde{H}=iH$ and $t=\beta$ Gao et al. (2024).
@@
We are grateful to Zongping Gong, Masaya Nakagawa, Shinsei Ryu, Yunfeng Jiang, Kohei Kawabata, and Masahito Ueda for fruitful discussion. H. L. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo.
Supplemental Material for
“Yang-Lee Zeros in Quantum Phase Transition: An Entanglement Perspective”
Appendix S1 Correlation Function in the SSH model
@@
Here we consider the Yang-Lee zeros in the non-Hermitian SSH model. The Bloch Hamiltonian of non-Hermitian SSH model is given by Chang et al. (2020)
$H_{k}=iu\sigma_{z}+(v+w\cos k)\sigma_{x}+w\sin k\sigma_{y}=\left(\begin{array}[]{cc}iu&we^{-ik}+v\\ we^{ik}+v&-iu\end{array}\right).$ (S.1)
@@
The energy spectrum is $E_{k,\pm}=\pm\sqrt{|we^{-ik}+v|^{2}-u^{2}}\equiv\pm E_{k}$. There are overall two phases in the Hamiltonian: the PT-symmetric phase for $|w-v|>u$ and the PT-broken phase for $|w-v|<u$. In the PT-symmetric phase, the Hamiltonian is gapped with the energy gap $\Delta=2\sqrt{(w-v)^{2}-u^{2}}$. In the PT-broken phase, the Hamiltonian is gapless with $\text{Re}[E_{k}]=0$ between the exceptional points given by
$k_{E}=\pm\arccos\frac{u^{2}-v^{2}-w^{2}}{2vw},$ (S.2)
@@
where the gap closes and the Hamiltonian cannot be diagonalized. The excitation spectrum around the exceptional points is given by $E_{k}\sim\sqrt{|k-k_{E}|}$ Chang et al. (2020).
@@
$H$ $=$ $\sum_{k}\left(\begin{array}[]{cc}c_{kA}^{\dagger}&c^{\dagger}_{kB}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{c}c_{kA}\\ c_{kB}\end{array}\right)$ (S.8)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&\sin\frac{\phi_{k}}{2}\\ -\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\\ \sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right)$ (S.18)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}E_{k}&0\\ 0&-E_{k}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right),$ (S.24)
@@
$\gamma_{kA}$ $=$ $\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}+\sin\frac{\phi_{k}}{2}c_{kB},$ (S.25)
$\gamma_{kB}$ $=$ $-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB},$ (S.26)
$\bar{\gamma}_{kA}$ $=$ $\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},$ (S.27)
$\bar{\gamma}_{kB}$ $=$ $-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB}.$ (S.28)
@@
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}A}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (S.29)
@@
$\{\bar{\gamma}_{kB},\gamma_{k^{\prime}B}\}$ $=$ $\left\{-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (S.30)
@@
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}B}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB}\right\}$ (S.31)
@@
Hence, we can still consider $\bar{\gamma}_{kA(B)}$ as the creation operator of the fermion with the flavor $A(B)$ and the momentum $k$, similarly with the Bogoliubov transformation in Yamamoto et al. (2019). Meanwhile, we can calculate the correlation functions with the linear transformations (S.25)-(S.28). Here we define the correlation function as
$C_{\alpha\beta}(x)=_{L}\langle c_{\alpha}^{\dagger}(x)c_{\beta}(0)\rangle_{R}:=\frac{1}{Z}\langle G_{L}|c_{\alpha}^{\dagger}(x)c_{\beta}(0)e^{-\beta E_{ks}}|G_{R}\rangle,$ (S.32)
@@
$C_{AA}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kA}\rangle_{R}$ (S.33)
@@
$C_{BB}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kB}\rangle_{R}$ (S.34)
@@
$C_{AB}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kB}\rangle_{R}$ (S.35)
@@
$C_{BA}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kA}\rangle_{R}$ (S.36)
@@
$C_{AA}(x)$ $=$ $\int\frac{dk}{2\pi}\langle c_{kA}^{\dagger}c_{kA}\rangle e^{ikx}$ (S.37)
@@
$C_{AB}(x)$ $=$ $\int\frac{dk}{2\pi}\left[\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\left(\frac{1}{e^{\beta E_{k}}+1}-\frac{1}{e^{-\beta E_{k}}+1}\right)\right]e^{ikx}$ (S.41)
@@
$\lim_{x\rightarrow\infty}C_{\alpha\beta}(x)\sim\frac{e^{-x/\xi}}{\sqrt{x}},$ (S.42)
which originates from the scaling behavior of $K_{0}(x)$. This exponenial decay is attributed to the finite energy gap of the system. From the definition of the critical exponents in the main text, we obtain $\nu=1,\eta=3/2$.
Appendix S2 Entanglement Entropy in Non-interacting Fermionic Systems
@@
$S_{A}=-\text{Tr}[\rho_{A}\log\rho_{A}],\rho_{A}:=\text{Tr}_{\bar{A}}\rho.$ (S.43)
The ground state here is defined as the eigenstate with only the negative-energy modes filled. We here consider the case of ground-state degeneracy where the system has two phases: the PT-broken phase and the PT-unbroken phase. The system has pure imaginary eigenvalues in the PT-broken phase which indicates that the phase is gapless and only finite real values in the PT-unbroken phase which indicates that the phase is gapped. In the PT-broken phase, the pure imaginary eigenvalues are surrounded by the exceptional points in the Brillouin zone, namely exceptional rings Budich et al. (2019); Bergholtz et al. (2021). The partition function defined in Eq. (5) in the main text can be written as
$Z=\prod_{k,\alpha}\left(1+e^{-\beta E_{k}^{\alpha}}\right),$ (S.44)
@@
$\text{Im}E_{k}^{\alpha}=\frac{(2n+1)\pi}{\beta}.$ (S.45)
We can infer that there always exist momenta $k$s satisfying Eq. (S.45) at absolute zero in the PT-unbroken phase, where Yang-Lee zeros emerge. Next, we are going to prove that the entanglement entropy shows at least a subarea-law scaling in the PT-unbroken phase. To show this, we follow the method used in Ref. Wolf (2006). The correlation matrix is defined as
$\gamma_{i\alpha,j\beta}=\delta_{ij}\delta_{\alpha\beta}-2\text{Tr}[\rho c^{\dagger}_{i\alpha}c_{j\beta}]=\langle\psi_{0}^{L}|(I_{ij}I_{\alpha\beta}-2c^{\dagger}_{i\alpha}c_{j\beta})|\psi_{0}^{R}\rangle,$ (S.46)
where $i,j\in\mathbb{Z}$ are the positions of the fermions in the subsystem $A$, $\alpha,\beta$ are the band indices and $I$ is the identity matrix. To simplify the calculation, we here assume the subsystem is a hypercube with edge length $L$. The entanglement entropy can be expressed with the eigenvalues $\lambda_{M}$ of the correlation matrix in the subsystem $A$ Chang et al. (2020).
$S_{A}$ $=\sum_{M=1}^{NL^{d}}h(\lambda_{M}),$ (S.47)
$h(x)$ $=-\frac{1+x}{2}\log\left(\frac{1+x}{2}\right)-\frac{1-x}{2}\log\left(\frac{1-x}{2}\right),$ (S.48)
@@
$\gamma_{i\alpha,j\beta}=\int d^{d}k(1-2\eta(k))e^{ik(i-j)}\delta_{\alpha\beta},$ (S.49)
@@
$S_{A}>\frac{1}{2}\text{Tr}[I-\gamma^{2}].$ (S.51)
With the introduction of the kernel Wolf (2006)
$F_{L}(x)=\sum_{i,j\in\mathbb{Z}_{L}}e^{ix(i-j)}=\frac{\cos(Lx)-1}{\cos(x)-1},$ (S.52)
@@
where $c_{-}:=\frac{2\sqrt{d}}{(2\pi)^{d+1}}s_{-}$. The entanglement entropy shows at least a logarithmic scaling when $d=1$. For the multi-band cases, we have $S_{A}>\lfloor N/2\rfloor c_{-}\log L$, where $\lfloor x\rfloor$ is the floor function. Hence, the PT-unbroken phase with Yang-Lee zeros is always accompanied by at least a subarea-law scaling for the entanglement entropy. In the PT-broken phase, the entanglement entropy exhibits area law since it is gapped. Hence, the edges of the distribution of Yang-Lee zeros in one-dimensional cases always correspond to the entanglement transition. For the other cases where the Fermi surface is not the exceptional ring, the discussion is exactly the same as the Hermitian one in Ref. Wolf (2006). Since Yang-Lee zeros can only take place in the gapless PT-broken phase, the edge of the Yang-Lee zeros can always correspond to the entanglement transitions.
Appendix S3 Yang-Lee Zeros in the Transverse-Field XY Model
For the transverse-field XY model, we find the Yang-Lee zeros by extending the magnetic field into a complex one. The model is given by
$H=J\sum_{i=1}^{L}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y})+h\sum_{i=1}^{L}\sigma_{i}^{z}.$ (S.58)
Under the Jordan-Wigner transformation Mbeng et al. (2020):
$\sigma_{i}^{x}=\prod_{l<i}(1-2c^{\dagger}_{l}c_{l})\frac{1}{2}(c_{i}+c_{i}^{\dagger}),\sigma_{i}^{y}=\prod_{l<i}(1-2c^{\dagger}_{l}c_{l})\frac{1}{2i}(c_{i}-c_{i}^{\dagger}),\sigma_{i}^{z}=c_{i}^{\dagger}c_{i}-\frac{1}{2},$ (S.59)
we obtain
$H$ $=$ $-\frac{J}{2}\sum_{i=1}^{L}(c_{i}^{\dagger}c_{i+1}+c_{i+1}^{\dagger}c_{i})+h\sum_{i=1}^{L}\left(c_{i}^{\dagger}c_{i}-\frac{1}{2}\right).$ (S.60)
Then we diagonalize it by Fourier transformation and get
$H=\sum_{k}(-J\cos k+h)c_{k}^{\dagger}c_{k}-\frac{hL}{2}=\sum_{k}E_{k}c_{k}^{\dagger}c_{k}-\frac{hL}{2},$ (S.61)
where $E_{k}=-J\cos k+h$. The partition function of the fermionic Hamiltonian is given by
$Z=\text{Tr}e^{-\beta H}=\prod_{k}(1+e^{-\beta E_{k}}).$ (S.62)
There are two phases: the gapless XY phase $h<J$ and the gapped paramagnetic phase $h>J$. To satisfy the condition of Yang-Lee zeros, we need a momentum $k$ such that
$1+e^{-\beta E_{k}}=0\Rightarrow-J\cos k+h=i(2n+1)\pi/\beta.$ (S.63)
If we generalize the magnetic field to a complex one, we find that this can be satisfied for $-J<\text{Re}[h]<J$, which corresponds to the gapless phase. Hence, the Yang-Lee zeros exist in the whole region of the gapless phase in the zero-temperature case.
Appendix S4 Derivation of Eqs. (16) and (17)
We here consider the 1D XXZ model with the Hamiltonian
$H=-J\sum_{i=1}^{L}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}+\Delta S_{i}^{z}S_{i+1}^{z}),$ (S.64)
@@
where $J>0$ and $\Delta$ is the anisotropy parameter. This model can be exactly solved by Bethe ansatz. The Bethe ansatz equations of the XXZ model are given by Takahashi (1999)
$\left(\frac{\sin\frac{\phi}{2}(x_{j}+i)}{\sin\frac{\phi}{2}(x_{j}-i)}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l}+2i)}{\sin\frac{\phi}{2}(x_{j}-x_{l}-2i)}.$ (S.65)

### v2

@@ IV Conclusion and Outlook
In this work, we have investigated the distribution of Yang-Lee zeros in quantum phase transitions of the SSH model and the XXZ model and shown how Yang-Lee zeros enter the phase transition points. In particular, in the XXZ model, we have proposed the simplified Bethe ansatz equation (SBAE) to exactly solve the energy levels around the ground state and calculated the partition function in the zero-temperature limit. We have also explored the critical behavior at the edges of Yang-Lee zeros of both models. Furthermore, we find that the edges of Yang-Lee zeros at absolute zero correspond to the entanglement transition points of the ground state and discuss the mechanism behind the correspondence. We have also generalized the correspondence to an arbitrary PT-symmetric free-fermionic system.
@@
The Yang-Lee zeros in quantum phase transitions studied in this paper are not only an interesting mathematical concept but also experimentally realizable. The non-Hermitian SSH model can be realized by dilating it into a larger Hermitian Hamiltonian due to the PT symmetry [46, 47, 48]. We can introduce an ancillary qubit as an environment coupled to the targeted spin to realize the Hamiltonian (2). For the non-Hermitian XXZ model, we can realize it in a non-Hermitian Bose-Hubbard model in a strongly correlated regime [44, 49, 50], where the non-Hermiticity originates from the two-body loss in the open quantum systems. By coupling the target system to an ancilla, we expect the Yang-Lee zeros and its associated edge singularity to be measurable with the decoherence of the spin in the ancilla [51, 52, 53, 19]. The Yang-Lee zeros in those systems can also be measured from the diagonal elements of the time-evolution operator $U=\exp(i\tilde{H}t)$, where $\tilde{H}=iH$ and $t=\beta$ [54]. To be specific, we can measure the partition function of a non-Hermitian SSH model by preparing a state whose dynamics is described by the $2\times 2$ Bloch Hamiltonian $iH_{k}$ in the momentum space, where $H_{k}$ is given in Eq. (2). Then the trace of the evolution operator gives the partition function: $Z=\prod_{k}\text{Tr}U_{k}$ with $t=\beta$ and $U_{k}=\exp{[i(iH_{k})t]}=\exp{(-H_{k}t)}$. Similarly, we can also prepare a state whose dynamics is governed by the Hamiltonian $iH$ with tomography [55] to measure the partition function of the non-Hermitian XXZ model where $H$ is shown in Eq. (9). The partition function is given by $Z=\text{Tr}U=\text{Tr}\exp{[i(iH)t]}=\text{Tr}\exp{(-Ht)}$ where $t=\beta$. The transition of the entanglement entropy is expected to be observed in the quantum gas microscope [56].
V ACKNOWLEDGMENTS
@@
We are grateful to Zongping Gong, Masaya Nakagawa, Shinsei Ryu, Yunfeng Jiang, Kohei Kawabata, and Masahito Ueda for fruitful discussion. H. L. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, at the University of Tokyo. This work was also supported by JSPS KAKENHI Grants No. 24KJ0824.
Appendix A Correlation Function in the SSH model
@@
Here we consider the Yang-Lee zeros in the non-Hermitian SSH model. The Bloch Hamiltonian of non-Hermitian SSH model is given by [35]
$H_{k}=iu\sigma_{z}+(v+w\cos k)\sigma_{x}+w\sin k\sigma_{y}=\left(\begin{array}[]{cc}iu&we^{-ik}+v\\ we^{ik}+v&-iu\end{array}\right).$ (20)
@@
The energy spectrum is $E_{k,\pm}=\pm\sqrt{|we^{-ik}+v|^{2}-u^{2}}\equiv\pm E_{k}$. Overall, there are two phases in the Hamiltonian: the PT-symmetric phase for $|w-v|>u$ and the PT-broken phase for $|w-v|<u$. In the PT-symmetric phase, the Hamiltonian is gapped with the energy gap $\Delta=2\sqrt{(w-v)^{2}-u^{2}}$. In the PT-broken phase, the Hamiltonian is gapless with $\text{Re}[E_{k}]=0$ between the exceptional points given by
$k_{E}=\pm\arccos\frac{u^{2}-v^{2}-w^{2}}{2vw},$ (21)
@@
where the gap closes and the Hamiltonian cannot be diagonalized. The excitation spectrum around the exceptional points is given by $E_{k}\sim\sqrt{|k-k_{E}|}$ [35].
@@
$H$ $=$ $\sum_{k}\left(\begin{array}[]{cc}c_{kA}^{\dagger}&c^{\dagger}_{kB}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{c}c_{kA}\\ c_{kB}\end{array}\right)$ (27)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&\sin\frac{\phi_{k}}{2}\\ -\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{cc}iu&v_{k}\\ v_{k}^{\ast}&-iu\end{array}\right)\left(\begin{array}[]{cc}\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}&-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}\\ \sin\frac{\phi_{k}}{2}&\cos\frac{\phi_{k}}{2}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right)$ (37)
$=$ $\sum_{k}\left(\begin{array}[]{cc}\bar{\gamma}_{kA}&\bar{\gamma}_{kB}\end{array}\right)\left(\begin{array}[]{cc}E_{k}&0\\ 0&-E_{k}\end{array}\right)\left(\begin{array}[]{c}\gamma_{kA}\\ \gamma_{kB}\end{array}\right),$ (43)
@@
$\gamma_{kA}$ $=$ $\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}+\sin\frac{\phi_{k}}{2}c_{kB},$ (44)
$\gamma_{kB}$ $=$ $-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB},$ (45)
$\bar{\gamma}_{kA}$ $=$ $\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},$ (46)
$\bar{\gamma}_{kB}$ $=$ $-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB}.$ (47)
@@
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}A}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (48)
@@
$\{\bar{\gamma}_{kB},\gamma_{k^{\prime}B}\}$ $=$ $\left\{-\frac{v_{k}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\cos\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k^{\prime}}^{\ast}}{|v_{k^{\prime}}|}\sin\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}A}+\cos\frac{\phi_{k^{\prime}}}{2}c_{k^{\prime}B}\right\}$ (49)
@@
$\{\bar{\gamma}_{kA},\gamma_{k^{\prime}B}\}$ $=$ $\left\{\frac{v_{k}}{|v_{k}|}\cos\frac{\phi_{k}}{2}c_{kA}^{\dagger}+\sin\frac{\phi_{k}}{2}c^{\dagger}_{kB},-\frac{v_{k}^{\ast}}{|v_{k}|}\sin\frac{\phi_{k}}{2}c_{kA}+\cos\frac{\phi_{k}}{2}c_{kB}\right\}$ (50)
@@
Hence, we can still consider $\bar{\gamma}_{kA(B)}$ as the creation operator of the fermion with the flavor $A(B)$ and the momentum $k$, similar to the non-Hermitian Bogoliubov transformation in Ref. [57]. Meanwhile, we can calculate the correlation functions with the linear transformations (44)-(47). Here we define the correlation function as
$C_{\alpha\beta}(x)=_{L}\langle c_{\alpha}^{\dagger}(x)c_{\beta}(0)\rangle_{R}:=\frac{1}{Z}\langle G_{L}|c_{\alpha}^{\dagger}(x)c_{\beta}(0)e^{-\beta E_{ks}}|G_{R}\rangle,$ (51)
@@
$C_{AA}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kA}\rangle_{R}$ (52)
@@
$C_{BB}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kB}\rangle_{R}$ (53)
@@
$C_{AB}(k)$ $=$ ${}_{L}\langle c_{kA}^{\dagger}c_{kB}\rangle_{R}$ (54)
@@
$C_{BA}(k)$ $=$ ${}_{L}\langle c_{kB}^{\dagger}c_{kA}\rangle_{R}$ (55)
@@
$C_{AA}(x)$ $=$ $\int\frac{dk}{2\pi}\langle c_{kA}^{\dagger}c_{kA}\rangle e^{ikx}$ (56)
@@
$C_{AB}(x)$ $=$ $\int\frac{dk}{2\pi}\left[\frac{v_{k}^{\ast}}{|v_{k}|}\cos\frac{\phi_{k}}{2}\sin\frac{\phi_{k}}{2}\left(\frac{1}{e^{\beta E_{k}}+1}-\frac{1}{e^{-\beta E_{k}}+1}\right)\right]e^{ikx}$ (60)
@@
$\lim_{x\rightarrow\infty}C_{\alpha\beta}(x)\sim\frac{e^{-x/\xi}}{\sqrt{x}},$ (61)
which originates from the scaling behavior of $K_{0}(x)$. This exponential decay is attributed to the finite energy gap of the system. From the definition of the critical exponents in the main text, we obtain $\nu=1,\eta=3/2$.
Appendix B Entanglement Entropy in Non-interacting Fermionic Systems
@@
$S_{A}=-\text{Tr}[\rho_{A}\log\rho_{A}],\rho_{A}:=\text{Tr}_{\bar{A}}\rho.$ (62)
The ground state here is defined as the eigenstate with only the negative-energy modes filled. Here we consider the case of ground-state degeneracy where the system has two phases: the PT-broken phase and the PT-unbroken phase. The system has pure imaginary eigenvalues in the PT-broken phase which indicates that the phase is gapless and only finite real values in the PT-unbroken phase which indicates that the phase is gapped. In the PT-broken phase, the pure imaginary eigenvalues are surrounded by the exceptional points in the Brillouin zone, namely exceptional rings [58, 59]. The partition function defined in Eq. (5) in the main text can be written as
$Z=\prod_{k,\alpha}\left(1+e^{-\beta E_{k}^{\alpha}}\right),$ (63)
@@
$\text{Im}E_{k}^{\alpha}=\frac{(2n+1)\pi}{\beta}.$ (64)
We can infer that there always exist momenta $k$s satisfying Eq. (64) at absolute zero in the PT-unbroken phase, where Yang-Lee zeros emerge. Next, we are going to prove that the entanglement entropy shows at least a subarea-law scaling in the PT-unbroken phase. To show this, we follow the method used in Ref. [40]. The correlation matrix is defined as
$\gamma_{i\alpha,j\beta}=\delta_{ij}\delta_{\alpha\beta}-2\text{Tr}[\rho c^{\dagger}_{i\alpha}c_{j\beta}]=\langle\psi_{0}^{L}|(I_{ij}I_{\alpha\beta}-2c^{\dagger}_{i\alpha}c_{j\beta})|\psi_{0}^{R}\rangle,$ (65)
where $i,j\in\mathbb{Z}$ are the positions of the fermions in the subsystem $A$, $\alpha,\beta$ are the band indices and $I$ is the identity matrix. To simplify the calculation, here we assume that the subsystem is a hypercube with edge length $L$. The entanglement entropy can be expressed with the eigenvalues $\lambda_{M}$ of the correlation matrix in the subsystem $A$ [35].
$S_{A}$ $=\sum_{M=1}^{NL^{d}}h(\lambda_{M}),$ (66)
$h(x)$ $=-\frac{1+x}{2}\log\left(\frac{1+x}{2}\right)-\frac{1-x}{2}\log\left(\frac{1-x}{2}\right),$ (67)
@@
$\gamma_{i\alpha,j\beta}=\int d^{d}k(1-2\eta(k))e^{ik(i-j)}\delta_{\alpha\beta},$ (68)
@@
$S_{A}>\frac{1}{2}\text{Tr}[I-\gamma^{2}].$ (70)
With the introduction of the kernel [40]
$F_{L}(x)=\sum_{i,j\in\mathbb{Z}_{L}}e^{ix(i-j)}=\frac{\cos(Lx)-1}{\cos(x)-1},$ (71)
@@
where $c_{-}:=\frac{2\sqrt{d}}{(2\pi)^{d+1}}s_{-}$. The entanglement entropy shows at least a logarithmic scaling when $d=1$. For the multi-band cases, we have $S_{A}>\lfloor N/2\rfloor c_{-}\log L$, where $\lfloor x\rfloor$ is the floor function. Hence, the PT-unbroken phase with Yang-Lee zeros is always accompanied by at least a subarea-law scaling for the entanglement entropy. In the PT-broken phase, the entanglement entropy exhibits area law since it is gapped. Hence, the edges of the distribution of Yang-Lee zeros in one-dimensional cases always correspond to the entanglement transition. For the other cases where the Fermi surface is not the exceptional ring, the discussion is exactly the same as the Hermitian one [40]. Since Yang-Lee zeros can only take place in the gapless PT-broken phase, the edge of the Yang-Lee zeros can always correspond to the entanglement transitions.
Appendix C Derivation of Eqs. (16) and (17)
We here consider the one-dimensional (1D) XXZ model with the Hamiltonian
@@
where $J>0$ and $\Delta$ is the anisotropy parameter. This model can be exactly solved by the Bethe ansatz. The Bethe ansatz equations of the XXZ model are given by [60]
$\left(\frac{\sin\frac{\phi}{2}(x_{j}+i)}{\sin\frac{\phi}{2}(x_{j}-i)}\right)^{L}=\prod_{l\neq j}\frac{\sin\frac{\phi}{2}(x_{j}-x_{l}+2i)}{\sin\frac{\phi}{2}(x_{j}-x_{l}-2i)}.$ (78)